OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the equation 15x - 5y = 30. Answer x-intercept = (-2, 0); y-intercept = (0, 6) x-intercept = (2, 0); y-intercept = (0, -6) x-intercept = (0, -6); y-intercept = (2, 0) x-intercept = (0, 6); y-intercept = (-2, 0)

OpenStudy (anonymous):

15x-5y = 30 -5y = 30-15x y = 3x-6 y = 3(0)-6 y = 0-6 y = -6, (0,6) (0) = 3x-6 0 = 3x-6 -3x = -6 x = 2, (2,0) Solution: x-intercept = (2,0); y-intercept = (0,-6)

OpenStudy (anonymous):

thanks